OpenStudy (anonymous):

2.34 as a percent and as a fraction?

OpenStudy (anonymous):

234%

OpenStudy (anonymous):

Percent = .0234 Fraction 2.34 / 100 = 234 / 10000 = (117/5000)

OpenStudy (anonymous):

Thank you!